### v1

@@ 2 Background: Large Language Models and Mechanistic Interpretability
Logit lens.
3 Preliminaries: Zero-shot Learning with MCQs
Zero-shot learning.
Table 1: The distribution of correct options on each training dataset. IOI, LD, and Greater-than datasets are manual-synthesised and we did not choose or place the correct option at A. For test datasets, we only select samples whose correct options are not ‘A’ to avoid overlap between anchored predictions from GPT2 models and the correct option. ()($\cdot$)( ⋅ ) indicates the number of choices.
Datasets Train Test A (%) B (%) C (%) D (%) E (%)
IOI (2) - 1000 0 100 - - -
LD (3) - 200 0 50 50 - -
Greater (4) - 1000 0 33.33 33.33 $3$
ARC (4) 1.12k 907 20.82 26.18 25.65 25.29 -
CSQA (5) 9.74k 982 19.60 20.25 19.98 20.38 19.79
Datasets and models.
@@ 3 Preliminaries: Zero-shot Learning with MCQs
To comprehensively verify and evaluate the anchored bias of GPT2 family, we consider 5 datasets, which include different numbers of choices from 2 to 5. Indirect Object Identification (IOI) Wang et al., (2022): IOI is a manually synthesised corpus used to understand a specific natural language task, where sentences such as “Afterwards Lisa and Rachel went to the garden, and Lisa gave a bone to” should be followed with “Rachel”. When two names are included in such sentences, the predicted name should not be the subject of the last clause. This task has been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024). However, we found that GPT2 family immediately fails this task if the input prompt is formatted as MCQ, where the incorrect subject of the last clause is placed in the ‘A’ choice. Greater-than task (Greater) Hanna et al., (2024): Greater-than task is also a manually synthesised corpus, which is used to evaluate GPT2’s ability to sentences such as “The war lasted from the year 1732 to the year 17”, and model will predict valid two-digit end years, i.e., years > 32. However, we found the same anchored bias like IOI when this task is formatted as MCQ in Fig. 1, and GPT2 family also fails to predict valid years and the prediction is always anchored at incorrect choice ‘A’. Logical Deduction of the Big-Bench (LD) [^3] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. ARC-Challenge (ARC) Clark et al., (2018) and CommensenseQA (CSQA) Talmor et al., (2019) are commonly-used MCQ benchmarks to evaluate LLMs. For models, we comprehensively evaluate the GPT2 family, i.e., GPT2-Small-124M, GPT2-Medium-355M, GPT2-Large-774M, and GPT2-XL-1.5B Radford et al., (2019).
Dist. (%) GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
IOI (2) 45.5 97.4 100.0 85.8
LD (3) 63.0 94.0 100.0 17.0
Greater (4) 32.1 95.0 99.5 98.0
ARC (4) 54.6 91.6 97.6 69.9
CSQA (5) 34.8 81.5 99.6 97.7
Evaluation metrics.
4 Discovering Anchored Bias in MCQs
Frequency of anchored bias in GPT2 family across all datasets.
As introduced in § 3, we use 5 different datasets with choices from 2 to 5 to investigate the anchored bias. Table 1 shows that the correct choices distribution of ARC and CSQA training dataset is balanced from ‘A’ to ‘E’. In addition, IOI, LD and Greater test datasets are manually synthesised, and we did not choose or place the correct choice at ‘A’. For all test datasets, we only select samples whose correct choice is not ‘A’ to avoid introducing extra bias, i.e., the mix-up between correct prediction and anchored bias of GPT2 family. Based on the randomly sampled test datasets in Table 1, we further calculate the distribution of anchored bias ‘A’ happened within each test dataset in Table 2. We can find that GPT2-Large and GPT2-Medium have the most serious anchored bias, and GPT2-XL and GPT2-Small have relatively less serious issues. Based on this situation, we mainly focus on investigating test samples which have anchored bias for each dataset. Table 7 in Appendix shows the number of test samples for each dataset and GPT2 model, where Infer. is used to localise and mitigate anchored bias and Eva. is used to verify the performance of mitigation.
Locating MLP of GPT2 family for anchored bias.
[…]
We calculate the MLP logit difference between anchored bias token ‘A’ and correct choice token B/C/D/E averaged across all layers and datasets for each GPT2 model using Infer. test samples. As shown in Fig. 3, layer 9 in GPT2-Small, layer 20 in GPT2-Medium, layer 34 in GPT2-Large, and layer 37/38/44 in GPT2-XL are dominant layers [^5] In this work, the layer number and head number start from 0. related to anchored bias. In addition, these layers are much closer to the final layer of GPT2, which agrees with Geva et al., (2022); Gurnee et al., (2023)’s finding that higher layers in GPT2 are relevant to semantic concepts or complicated tasks. We also notice that the last one or two layers in each GPT2 model do not have anchored bias at all, and they contribute more logits to correct choice token ‘B/C/D/E’ to ‘A’.
Table 3 Vector column demonstrates the specific value vectors $\mathbf{v}^{\ell,n}$ which are responsible for anchored bias. For each value vector, we also calculate the percentage of how frequently it is recognised as an anchored-bias vector across all datasets for each GPT2 model. We can find that most identified value vectors have more than 50% chance with anchored bias happening across all datasets and different GPT2 models. To further verify whether these value vectors store anchored knowledge bias, we unembeded each value vector logit and selected the top-10 tokens with the highest probability. As shown in Table 3, we can find that most top-10 tokens within each value vector are relevant to ‘A’, e.g., ␣A, A, ␣a, a, ␣first, etc. This finding proves that some value vectors in $W_{\text{out}}$ of GPT2 family store knowledge bias after pertaining, and these knowledge biases will become anchored bias when the input prompt is formatted as MCQ. In addition, most unembeded tokens are stopwords, e.g., pronouns, articles, prepositions, etc, which also agrees with the findings that “stopwords/punctuation” are commonly distributed in the value vectors of MLP (Geva et al.,, 2022).
@@ 4 Discovering Anchored Bias in MCQs
[…]
$\mathbf{v}^{9,788}$ (51.0%) ␣and, ␣in, ␣to, ␣or, ␣a, ␣at, ␣the, ␣..., ␣that, "
$\mathbf{v}^{20,1731}(38.3\%)$first, ␣First, first, ␣FIRST, First, ␣1, 1, ␣Firstly, Firstly, ␣begin
$\mathbf{v}^{34,2103}(70.2\%)$ romeda, maxwell, {CJK}UTF8gothドラゴン, lvl, mobi, elaide, amsung, nil, 911, dylib
$\mathbf{v}^{34,4178}$ (54.7%) reply, interstitial, ␣emer, 76561, ␣err, sg, ␣whe, eers, oi, ␣ignor
$\mathbf{v}^{44,2995}$ (44.7%) ggles, atchewan, ␣aw, ␣let, eed, ␣ont, wn, ␣be, gg, ␣palp
$\mathbf{v}^{44,128}$ (11.9%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
$\mathbf{v}^{38,4191}$ (100%) ␣a, ␣an, a, ␣of, ,, ␣and, ., ␣in, an, ␣the
@@
$\mathbf{v}^{37,2966}$ (67.2%) ␣a, ␣an, ,, ␣and, ␣to, ␣the, ␣in, ␣(, ␣one, .
Figure 4: Attention pattern logit lens of the difference between anchored bias token ‘A’ and correct tokens (one of B,C,D,E), i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$ which is averaged within GPT-2 family across all layers and all datasets. The deeper the blue blocks are at each layer, the more serious the anchored bias is, and vice versa.
Locating attention heads of GPT2 family for anchored bias.
As explained in § 2, attention pattern $\mathbf{r}_{i,j}^{\ell,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$. We use logit lens to analyse the logit difference of final input prompt token contribution between anchored bias ‘A’ and correct choices ‘B/C/D/E’, i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$. As shown in Fig. 4, L8H1 and L10H8 in GPT2-Small, L18H12 and L20H5 in GPT2-Medium, L23H8 and L30H0 in GPT2-Large, L31H9 and L34H14666L34H14 indicates layer 34 and head 14 and both of them start from 0. in GPT2-XL are dominant heads related to anchored bias. Those heads are also distributed closer to the final layer in each GPT2 model. We zoom in on the L8H1 and L10H8 attention pattern of the final input token in GPT2-Small using a sample from IOI dataset. As shown in Fig. 5, the final token ‘:’ attends more weights on the anchored bias token ‘A’ than the correct choice token ‘B’, which agrees with our identified attention head using logit difference.
5 Mitigating Anchored Bias
Mitigating anchored bias in MLP.
According to the findings in § 4, we localise the specific value vector in MLP related to the anchored bias. We further aim to solve the following research question: Can we fix the identified value vector in MLP by updating its values?
[…]
@@ 5 Mitigating Anchored Bias
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) ␣B, B, ␣b, ␣C, ␣D, ␣P, ␣L, ␣R, ␣H, ␣F
$\mathbf{v}^{9,2859}$ (61.8%) ␣B, B, ␣b, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣F
$\mathbf{v}^{9,788}$ (51.0%) ␣B, ␣b, B, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣BC
$\mathbf{v}^{20,1731}(38.3\%)$C, C, ␣B, ␣c, ␣D, ␣CS, ␣F, ␣P, ␣G, ␣T
$\mathbf{v}^{34,2103}(70.2\%)$C, ␣c, C, ␣B, ␣D, ␣F, ␣P, ␣G, ␣T, ␣M
$\mathbf{v}^{34,4178}$ (54.7%) ␣C, ␣c, C, ␣B, ␣P, ␣D, ␣F, ␣G, ␣T, ␣L
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) ␣C, ␣c, C, ␣A, ␣B, ␣D, ␣F, ␣P, ␣T, ␣G
$\mathbf{v}^{44,2995}$ (44.7%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣T, ␣F, ␣R, ␣G
$\mathbf{v}^{44,128}$ (11.9%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
$\mathbf{v}^{38,4191}$ (100%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣L, ␣R
@@
$\mathbf{v}^{37,2966}$ (67.2%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣G, ␣L
where $\lambda_{1}=1,\lambda_{2}=8$. After updating the corresponding value vector in MLP, we utilise the updated GPT2 model to predict the next token of the same input MCQ prompts. We comprehensively evaluate each updated GPT2 model with the corresponding modified value vector using Infer. and Eva. across all datasets. As shown in Table 5, most updated value vectors achieve high classification accuracy regarding MCQ tasks, even with multiple near 100% or 100% accuracy in $\mathbf{v}^{9,1853}$ and $\mathbf{v}^{9,788}$ of GPT2-Small, $\mathbf{v}^{34,1541}$ of GPT2-Large, $\mathbf{v}^{44,4967}$ and $\mathbf{v}^{38,4191}$ of GPT2-XL. For those value vectors with around a 60-70% chance of anchored bias happening across all datasets and different GPT2 models (i.e., $\mathbf{v}^{9,2859}$ of GPT2-Small, $\mathbf{v}^{20,3713}$ of GPT2-Medium, $\mathbf{v}^{34,2103}$ of GPT2-Large, and $\mathbf{v}^{37,2966}$ of GPT2-XL), the classification accuracy still 68.09% averaged all datasets and all models. This means that the simple and straightforward method (i.e., Eq. 10) is effective, and we do not need to fine-tune the whole GPT2 model to fix the anchored bias. We further unembeded the updated value vectors in Table 4, and it shows that the anchored bias token ‘A’ is significantly removed and the new top-10 tokens for each new value vector are replaced with the correct choices token ‘B/C/D/E’.
Table 5: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model.

### v2

@@ Appendix C Detailed Explanations of LLMs Architecture
Appendix D Details of Each Dataset and Experimental Settings
@@
  • Indirect Object Identification (IOI) Wang et al., (2022): IOI is a manually synthesised corpus used to understand a specific natural language task, where sentences such as “Afterwards Lisa and Rachel went to the garden, and Lisa gave a bone to” should be followed with “Rachel”. When two names are included in such sentences, the predicted name should not be the subject of the last clause. This task has been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024). However, we found that GPT2 family immediately fails this task if the input prompt is formatted as MCQ, where the incorrect subject of the last clause is placed in the ‘A’ choice.
  • Greater-than task (Greater) Hanna et al., (2024): Greater-than task is also a manually synthesised corpus, which is used to evaluate GPT2’s ability to sentences such as “The war lasted from the year 1732 to the year 17”, and model will predict valid two-digit end years, i.e., years > 32. However, we found the same anchored bias like IOI when this task is formatted as MCQ in Fig. 1, and GPT2 family also fails to predict valid years and the prediction is always anchored at incorrect choice ‘A’.
  • Logical Deduction of the Big-Bench (LD) [^8] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. Each MCQ in the task includes three similar objects in a naturally ordered context (e.g., books of various colours sitting on a shelf) and a set of simple clues regarding their placement (e.g., "the red book is to the right of the green book") such that no clue is redundant. The challenge is to assign the highest probability to correct MCQ choice about which object lies at which position.
  • ARC-Challenge (ARC) Clark et al., (2018): ARC is a real grade-school level, multiple-choice science questions, which includes Easy and Challenge versions. We choose the Challenge version to evaluate the anchored bias.
  • CommensenseQA (CSQA) Talmor et al., (2019): CSQA is a new multiple-choice question answering dataset that requires different types of commonsense knowledge to predict the correct answers.
All GPT2 models are run during inference time and parameters inside of each GPT2 model are frozen. All experiments can be easily run using CPU or GPU, e.g., Apple Macbook Pro with M1 Pro chip or NVIDIA 3090 Ti with 24GB GPU RAM.
Appendix E The Statistic Information of Each Test Dataset for GPT2 Models
We split each dataset into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set for modified GPT2 model verification on the MCQ task. Table 7 shows the number of test data samples for each dataset.
Num. GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
IOI (2) 410 45 877 97 900 100 773 85
LD (3) 114 12 170 18 180 20 31 3
Greater (4) 289 32 856 95 897 99 883 98
ARC (4) 446 49 748 83 797 88 571 63
CSQA (5) 308 34 720 80 881 97 864 95
Appendix F Top-10 Unembeded Tokens from Identified Value Vectors
@@
GPT2-Small $\mathbf{v}^{9,788}$ (50.0%) ␣and, ␣in, ␣to, ␣or, ␣a, ␣at, ␣the, ␣..., ␣that, "
GPT2-Medium $\mathbf{v}^{20,1731}(38.2\%)$first, ␣First, first, ␣FIRST, First, ␣1, 1, ␣Firstly, Firstly, ␣begin
GPT2-Large $\mathbf{v}^{34,2103}(70.1\%)$ romeda, maxwell, {CJK}UTF8gothドラゴン, lvl, mobi, elaide, amsung, nil, 911, dylib
$\mathbf{v}^{34,4178}$ (55.4%) reply, interstitial, ␣emer, 76561, ␣err, sg, ␣whe, eers, oi, ␣ignor
GPT2-XL $\mathbf{v}^{44,2995}$ (27.5%) ggles, atchewan, ␣aw, ␣let, eed, ␣ont, wn, ␣be, gg, ␣palp
$\mathbf{v}^{44,128}$ (22.7%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
@@
$\mathbf{v}^{37,2966}$ (66.2%) ␣a, ␣an, ,, ␣and, ␣to, ␣the, ␣in, ␣(, ␣one, .
We unembedded more identified anchored-bias value vectors in Table 8, and we can find there are several tokens related to ‘A’, such as ␣a, ␣first, ␣First, ␣1, ␣A.
Appendix G Top-10 Unembedded Tokens from Updated Value Vectors
@@
GPT2-Small $\mathbf{v}^{9,788}$ (50.0%) ␣B, ␣b, B, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣BC
GPT2-Medium $\mathbf{v}^{20,1731}$ (38.2%) ␣C, C, ␣B, ␣c, ␣D, ␣CS, ␣F, ␣P, ␣G, ␣T
GPT2-Large $\mathbf{v}^{34,2103}$ (70.1%) ␣C, ␣c, C, ␣B, ␣D, ␣F, ␣P, ␣G, ␣T, ␣M
$\mathbf{v}^{34,4178}$ (55.4%) ␣C, ␣c, C, ␣B, ␣P, ␣D, ␣F, ␣G, ␣T, ␣L
GPT2-XL $\mathbf{v}^{44,2995}$ (27.5%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣T, ␣F, ␣R, ␣G
$\mathbf{v}^{44,128}$ (22.7%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
@@
$\mathbf{v}^{37,2966}$ (66.2%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣G, ␣L
After directly updating each identified value vector from MLP, we further unembedded them and selected the top-10 tokens with the highest probability. Table 9 shows that those top-10 tokens changes from original ‘A’ to other correct choice tokens, i.e., B, C, D, E.
Appendix H Ablation Study of Different $\lambda_{2}$ in Eq. 8
We conduct an ablation study by changing $\lambda_{2}$ from 8 to 2 in Eq. 8, and evaluate the classification accuracy of each Infer. and Eva. dataset using GPT2 family. Table 10,11,12,13,14,15,16 show the ablation study results under different $\lambda_{2}$. We can find that the classification accuracy shows a decreasing trend with $\lambda_{2}$ from 8 to 2, which indicates that the effect of direct value vector update using Eq. 8 decreases.
Table 10: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=8$.